A bullet weighing 4.5*10^-3 kg is fixed horizontally into a 1.8 kg wooden block at rest on a horizontal surface. The coefficient of kinetic friction between block and surface is 0.20. The bullet comes to rest in the block which moves 1.8m. Find the speed of the bullet.
1
Expert's answer
2021-09-08T08:14:11-0400

Given:

"m=4.5*10^{-3}\:\rm kg"

"M=1.8\:\rm kg"

"\mu=0.20"

"d=1.8\:\rm m"


Let the initial speed of the bullet we denoted as "u".Then initial speed of the block is given by the law of conservation of momentum

"v=u\frac{m}{M+m}\approx u\frac{m}{M}"

The energy-work theorem gives

"\frac{Mv^2}{2}=F_fd=\mu Mgd"

So,

"v=\sqrt{2\mu g d}=u\frac{m}{M}"

"u=\sqrt{2\mu g d}\frac{M}{m}"

"u=\sqrt{2*0.20*9.8*1.8}*\frac{1.8}{4.5*10^{-3}}\\=1063\:\rm m/s"